# Interpolative separable density fitting on adaptive real space grids

### v1

@@ III.5 Correlated electron structure using GWGW
We next assess the performance of our framework in correlated electronic structure calculations within the $GW$ approximation, using the THC decomposition of the ERI tensor introduced in Eqs. 4 and 5. Implementation details of the THC–$GW$ algorithm are described in Ref. 45.
The $GW$ self-energy consists of two components: the static Hartree-Fock potential (the Fock matrix $F$), which provides the mean-field description of electrons, and the dynamic self-energy $\Sigma(\tau)$, a function of imaginary time $\tau$, which captures dynamic electron correlations. As shown at the beginning of Sec. II, both terms depend explicitly on the ERI tensor, and are therefore directly influenced by the accuracy of the ISDF approximation.
@@
The ISDF errors in the Fock matrix $F$ are noticeably larger than those in the ERI tensor $V$ reported in Fig. 4, with chemical accuracy achieved only when $\alpha\approx 18$ or larger. This is consistent with prior observations in THC-based electronic structure methods, for which more auxiliary functions are needed to reach satisfactory HF accuracy, limiting the use of THC for HF calculations. More specifically, our data suggest that the dominant contribution to the Fock matrix error is the Hartree term, whereas the exchange potential shows convergence comparable to or better than that of the ERI tensor itself. This is likely a consequence of the larger overall magnitude of the Hartree potential. This result motivates a practical hybrid scheme for HF calculations: compute the Hartree potential directly using the adaptive grid DMK method and use ISDF only for the exchange term. The total cost of the former scales as $\mathcal{O}(MN^{2})$, less than the $\mathcal{O}(MR^{2})$ cost of our ISDF scheme. With this hybrid approach, chemical accuracy in the Fock matrix is attainable with $\alpha\approx 10$.
Fig. 7 also shows the maximum error of the dynamic $GW$ self-energy (5) across all orbitals and imaginary times. We observe convergence similar to that for the exchange potential, with $\alpha\approx 12$ sufficient to achieve chemical accuracy. Importantly, computing the $GW$ self-energy involves contractions over the full $V_{ijkl}$ tensor, including the occupied-virtual and virtual-virtual blocks. In contrast, the Hartree-Fock potential only depends on the occupied subspace. The fact that both the dynamic self-energy and the exchange potential exhibit consistent convergence is a consequence of the accurate representation of the full ERI tensor by the ISDF approach, confirming the robustness and transferability of the ISDF approximation beyond mean-field theory.
@@ IV Conclusion
Our approach addresses the primary computational bottleneck in THC-HF and THC-$GW$ calculations involving sharply localized basis functions, yielding a fully cubic-scaling workflow that makes large-scale all-electron calculations practical. The explicit inclusion of core orbitals is necessary for describing core-level excitations and for accurately capturing many-body renormalization effects between core and valence bands. We note that our framework can be straightforwardly generalized to periodic systems, for which the simultaneous separation of $k$-points and orbital indices makes ISDF even more advantageous. This only requires a modification of the underlying adaptive Poisson solver, and is a topic of our future work.

### v2

@@ III.5 Correlated electron structure using GWGW
Figure 8: Convergence of orbital energies and total energies with respect to the compression factor $\alpha$ for $(\mathrm{NH}_{3})_{2}$. The left panel shows the HOMO (solid lines) and LUMO (dashed lines) energies, while the right panel reports the correlation energy (solid lines) and the Hartree-Fock energy (dashed lines).
We next assess the performance of our framework in correlated electronic structure calculations within the $GW$ approximation, using the THC decomposition of the ERI tensor introduced in Eqs. 4 and 5. Implementation details of the THC–$GW$ algorithm are described in Ref. 46.
The $GW$ self-energy consists of two components: the static Hartree-Fock potential (the Fock matrix $F$), which provides the mean-field description of electrons, and the dynamic self-energy $\Sigma(i\omega)$, a function of Matsubara frequencies $i\omega$, which captures dynamic electron correlations. As shown at the beginning of Sec. II, both terms depend explicitly on the ERI tensor, and are therefore directly influenced by the accuracy of the ISDF approximation.
@@
The ISDF errors in the Fock matrix $F$ are noticeably larger than those in the ERI tensor $V$ reported in Fig. 4. This is consistent with prior observations in THC-based electronic structure methods, for which more auxiliary functions are needed to reach satisfactory HF accuracy, limiting the use of THC for HF calculations [37, 78]. More specifically, our data suggest that the dominant contribution to the Fock matrix error is the Hartree term, whereas the exchange potential shows convergence comparable to or better than that of the ERI tensor itself. This is likely a consequence of the larger overall magnitude of the Hartree potential. This result motivates a practical hybrid scheme for HF calculations: compute the Hartree potential directly using the adaptive grid DMK method and use ISDF only for the exchange term. The total cost of the former scales as $\mathcal{O}(MN^{2})$, less than the $\mathcal{O}(MR^{2})$ cost of our ISDF scheme. Within this hybrid approach, one could potentially reduce the required ISDF compression factor $\alpha$ even in an HF calculation.
Fig. 7 also shows the maximum error of the dynamic $GW$ self-energy in Eq. 5 across all orbitals and Matsubara frequencies. We observe convergence similar to that for the exchange potential. Importantly, computing the $GW$ self-energy involves contractions over the full $V_{ijkl}$ tensor, including the occupied-virtual and virtual-virtual blocks. In contrast, the Hartree-Fock potential only depends on the occupied subspace. The fact that both the dynamic self-energy and the exchange potential exhibit consistent convergence is a consequence of the accurate representation of the full ERI tensor by the ISDF approach, confirming the robustness and transferability of the ISDF approximation beyond mean-field theory.
To further assess the impact of ISDF errors on physical observables, Fig. 8 shows the convergence of orbital energies (left panel) and total energies (right panel) with respect to $\alpha$. Here the hybrid approach is adopted in which only the exchange and dynamic self-energy terms are evaluated with THC-ERIs.
The convergence of both the highest occupied molecular orbital (HOMO) and lowest unoccupied molecular orbital (LUMO) energies closely follow that of the exchange potential and dynamic self-energy, with chemical accuracy reached at $\alpha\approx 12$. Similar behavior is observed for $GW$ correlation eneries. Notably, the errors in HOMO/LUMO energies and correlation energies are consistently much smaller than the errors of the individual ERI elements (see Fig. 4). This suggests that the ERI components most relevant for correlation effects, which are dominated by valence orbitals, are well represented within the all-electron THC-ERI framework. By contrast, the HF energy contains large contributions from core orbitals, which are more sensitive to the THC-ERI errors, and therefore converges more slowly. While the hybrid approach significantly improves the accuracy of the HF energies, convergence still requires $\alpha\approx 16$. These results indicate that for total energy calculations requiring high precision, ISDF-compressed ERIs are most efficient for post-HF treatment.
@@ IV Conclusion
Our approach addresses the primary computational bottleneck in THC-HF and THC-$GW$ calculations involving sharply localized basis functions, yielding a fully cubic-scaling workflow that makes large-scale all-electron calculations practical. The explicit inclusion of core orbitals is necessary for describing core-level excitations and for accurately capturing many-body renormalization effects between core and valence bands.
An important direction for our future work is extending our framework to periodic systems, for which the simultaneous separation of $k$ point and orbital indices makes ISDF even more advantageous. This will involve generalizing the DMK algorithm to periodic boundary conditions for $\Gamma$-point calculations, as described in Ref. 79, and quasi-periodic boundary conditions for $k$-point sampling.